# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 5 Solutions for Preserving Privacy in LLMs: Model
Figure 6: Differential Privacy Applied on LMs. Differential privacy represents one of the main solutions used to preserve privacy when training a language model. We have already discussed Differential Privacy (DP) in Section 2.1; here, we contextualize DP with its usage in modern large language models that, most of the time, consist of adding noise in the training gradients (DP-SGD). On the right: Adding noise to the gradients of all the training parameters in a language model is often expensive. Parameter-efficient fine-tuning techniques can reduce the number of training parameters, thus lowering the cost of applying Differential Privacy to these gradients. Picture adapted from LoRA (Hu et al., 2021), a popular PEFT method.
In this section, we explore the various contemporary methods employed to ensure privacy preservation within machine learning models. The focus is on four primary approaches that operate directly on the model. Firstly, we explore the concept of Differential Privacy applied during the training phase (§5.1). This method involves training a model with techniques designed to protect the privacy of the data used by injecting noise (for an overview on Differential Privacy see §2.1). Secondly, we examine the application of Differential Privacy during the inference phase (§5.2, where privacy-preserving mechanisms are applied when the model is making predictions. Thirdly, we discuss federated learning with Differential Privacy (§5.3), which combines federated learning techniques with Differential Privacy to provide data safety in decentralized environments. Lastly, we discuss machine unlearning (§5.4), a technique that enables an existing model to forget or unlearn specific privacy-sensitive information. Each of these methods offers unique benefits and challenges in the pursuit of robust privacy protection in large language models.
@@ 5.1 Training Large Language Models with Differential Privacy
Training a Large Language Model (LLM) typically involves three sequential phases: pre-training on a broad collection of unannotated corpora (phase 1), fine-tuning on a specialized corpus using instruction-tuning procedures (phase 2), and final alignment according to human preferences (phase 3). We refer the reader to §2 for an in-depth overview of how these phases work. In this section, we propose methodologies that guarantee privacy by using differential privacy during the pre-training phase (phase 1 - §5.1.1) or apply it during the fine-tuning phase (phase 2 - §5.1.2). We also analyze methods that focus on Parameter Efficient Fine-Tuning techniques (§5.1.3) while leaving out of the scope of this survey works that focus on alignment via Reinforcement Learning and differential privacy since, to the best of our knowledge, there are very few works that just started exploring this aspect (Wu et al., 2024).
Despite the significant concerns over privacy, most of today’s LLMs abstain from any privacy-specific measure, relying instead on publicly available internet data for pre-training. While this data is public, it often contains sensitive information, posing risks that might harm individuals and undermine trust. Public data, thus, carries inherent risks, yet remains a widely used resource. On the other hand, fine-tuning often requires using private datasets that entities prefer to keep confidential to protect sensitive information or maintain a competitive edge. We refer to "public training" when the training process involves using datasets that are available to the general public and to "private training" when the training process relies on restricted, confidential data that is not available to the public and must be carefully protected for privacy reasons. Here, Differential Privacy could offer a solution, minimizing risks associated with data memorization and safeguarding the privacy of individuals in these datasets.
We have already explained Differential Privacy (DP) in Section 2.1; here, we contextualize DP with its usage in modern large language models. Specifically, working with DP often involves deciding where to inject noise into the training process, typically by adding noise to the gradients as shown in Step 2 of Figure 6. This approach generally involves using DP-SGD (Song et al., 2013), a variation of the classic SGD that clips gradient norms and then adds Gaussian noise to the average clipped gradient norm (see §2.1 for more details).
@@ 5.1.1 Pre-training with Differential Privacy
Pre-training a Large Language Model from scratch is usually a very cost-intensive procedure that just a few companies and institutions in the world can afford. In addition to this, adding differential privacy to this pre-training phase further increases the difficulties since: (i) adding noise increases the already exorbitant training cost, slowing the convergence time, and (ii) the training phase requires a vast collection of private data, which is often scarce and difficult to obtain. Thus, most models pre-trained from scratch with Differential Privacy are relatively small compared to current Large Language Model standards. Nevertheless, they offer both theoretical and practical insights into why fully private pre-training is considered a sub-optimal strategy. Instead, a mixed solution involving pre-training on public data is recommended. In this section, we will explain why. All the works discussed are also summarized in Table 4.
@@
It is crucial to acknowledge that the discussed works were capable of conducting full training from scratch largely because they utilized BERT (Devlin et al., 2019), which possesses a relatively modest 109M parameters, a figure significantly smaller than those of modern LLMs such as GPT-3 with 175B parameters (Brown et al., 2020b). Currently, the standard practice, even in settings where data are not private and there is no need to perform DP tuning, involves leveraging a pre-trained model and fine-tuning it for specific tasks. This approach not only reduces computational demands but also allows the model to harness a broad and diverse initial training dataset, enabling it to subsequently adapt more effectively to a private dataset and its specific domain, often enhancing performance. However, the benefits extend beyond performance improvements. Recent studies, including Kerrigan et al. (2020) that compare public pre-training with differentially private fine-tuning against entirely private DP training, demonstrate significantly better results with the former method. This provides empirical evidence that DP learning greatly benefits from an initial phase of public pre-training. Full training solely on a private dataset often results in high perplexity values, indicating a model that would likely be ineffective. In Kerrigan et al. (2020), the comparison of models fully trained privately versus those benefiting from public pre-training followed by private fine-tuning demonstrates that public pre-training is essential for enabling effective private learning. Additionally, since DP-SGD introduces noise during training, limiting this noise to the fine-tuning phase rather than the entire training process results in less overall noise and better performance. This improvement is due to the general (clean) knowledge acquired during the pre-training phase.
@@
Ganesh et al. (2023) present a study that compares the performance obtained via a public pre-training followed by private fine-tuning with a scenario where both stages are conducted in a differentially private manner. This paper emphasizes the theoretical underpinnings of Differential Privacy in training models, arguing that for a differentially private model to be effective on the task it is trained on, it is essential to first undergo pre-training on public data. The theoretical basis provided in the study explains that optimization involves two phases: initially, the model transitions from a random initialization towards a favourable basin, and subsequently, it addresses a convex optimization problem to find the local minimum within that basin. If attempted entirely within a private context, incorporating private data and noise addition, the initial stage would necessitate substantially more data to achieve comparable outcomes to those achievable with public data, as supported by experimental evidence. Moreover, given the typical scarcity of private data, it might be impractical or even impossible to gather sufficient data to effectively train a model privately from scratch. Therefore, pre-training on a larger public dataset not only aids in overcoming these limitations but also enhances the model’s resilience to distribution shifts between the public and private datasets.
@@
Tramèr & Boneh (2020) offer another perspective, arriving at a similar conclusion: public pre-training is crucial for achieving desirable outcomes in a private learning environment. Although this study is centered on image classification using CNNs, the findings reinforce that the advantages of public pre-training transcend various contexts, including this one. In their approach, the researchers enhance the performances of fully private training by incorporating handcrafted features, which yield improved results. Nevertheless, these outcomes still do not match those achieved through a combination of public pre-training and private fine-tuning. The collective insight from numerous studies is that a public pre-training phase is indispensable in private learning scenarios.
@@
While the works reviewed so far performed a "classical pre-training" (i.e., gaining knowledge from a large dataset), Yu et al. (2023) demonstrate that pre-training can also be conducted selectively, i.e., by focusing on a specific (smaller) subset of data that allows DP training to become more efficient. Their approach involves a novel technique of selecting a subset of the pre-training dataset, OpenWebText (Gokaslan & Cohen, 2019), whose distribution closely resembles the fine-tuning dataset’s distribution, Enron email dataset (Klimt & Yang, 2004). To filter the pre-training dataset, a classifier trained with Differential Privacy predicts whether a pre-training sample belongs to the distribution of the fine-tuning (private) dataset. Each sample in the pre-training dataset is then scored for how likely it is to belong to the fine-tuning dataset, then the top-k samples are used for public pre-training. This is followed by differentially private fine-tuning on the private dataset, with the final privacy loss taking into account both the classifier and the ultimate model. This method of dataset selection is designed to minimize distribution shifts, enabling smaller models to achieve better results. The results demonstrate strong performances, surpassing more complex baselines and thus enhancing efficiency to the point where this can be considered a model compression tool. However, a potential oversight of this method is the non-consideration of biases within the fine-tuning dataset, which could lead to the inadvertent transfer of these biases into the model through the selected pre-training samples. This is not as likely when the pre-training dataset is fully utilized. inline]While the learning process shows great benefit from a minimized distribution shift, this is not necessarily a good thing. As highlighted by Tramèr et al. (2024), this can make the results less meaningful, as it becomes unclear if we are looking at an actual accomplishment for private learning or just another example of intra-domain transfer learning. They argue that most literature work use for pre-training and fine-tuning two datasets that are too similar, leading to these results only being valid if the distributions are close enough and making them invalid for fields where the publicly available data distribution is different than the private data distribution (which is often the case in the medical field). They also argue that this is fine if we are in the situation (called jokingly transfermania) where this is the norm, but if in the real world the distributions are usually different (privacyland) these results are not really meaningful.
@@ 5.1.2 Fine-tuning with Differential Privacy
Pre-training a large language model from scratch, with or without Differential Privacy, has become wildly expensive given the sheer size of these models. Currently, the prevalent approach consists of using an already available LLM pretrained checkpoint and fine-tuning it for specific downstream tasks, including applications of Differential Privacy (DP). In this section, we survey several different works that apply Differential Privacy on top of already pre-trained LMs. These works are also available in tabular format in Table 5. Yue et al. (2022) show that using DP-SGD it is possible to successfully train LMs for text generation. They apply DP-SGD to fine-tune various sizes of GPT2, using the Open Yelp Dataset. This is mainly done to generate synthetic reviews that cannot be traced to the actual reviewer. The generated reviews are then used to train classifiers to predict ratings and business categories. The quality of the synthetic text is assessed by comparing its distribution to the original dataset (through F1-score, Fréchet Inception Distance Heusel et al. (2018), and MAUVE Pillutla et al. (2021)) and the classifier performance against models trained on the original data. In this task, the largest GPT2 models produce data that more closely resemble the original data, probably because they are less influenced by DP. Experiments indicate that DP introduces a minimal performance drop, often less than 0.01%, within a privacy budget of $\epsilon=4$. Remarkably, the GPT2-Large model with DP nearly matches the performance of its non-private counterpart, with only a 0.008% difference in accuracy on downstream tasks. Privacy testing was performed following the method by Carlini et al. (2019): this consists of injecting canaries inside the training set in order to extract them later on, to prove that the model is capable (or not) of memorizing training data. When these canaries were injected into the training data, the model could not be prompted to reproduce them under DP conditions, even with a repetition rate of 100. In contrast, without DP, the canaries were reproduced 80% of the time, demonstrating effective privacy preservation. This work highlights the feasibility of DP fine-tuning.
@@
When we investigate DP learning and the used techniques, there is a constant comparison with standard learning, as it is the most widely known and adopted. Papernot et al. (2019) specifically focus on the differences between private and public training, trying to shed some light on the best conditions for private training. They investigate architectures, parameter initializations and optimizers. We are not very interested in architectures as they do not consider transformers but only convolutional architectures and we already know from §5.1.1 that it’s best to start from parameters that were previously optimized through a public pre-trained. The novelty that this research points out is that adaptive optimizers like Adam provide marginal gains (if any) in this context. Putting in direct comparison DP-SGD and DP-Adam, they find that while the latter converges faster initially, it eventually slows down to the point where DP-SGD can outperform it, reaching a lower loss. Another relevant difference they point out is how different a working learning rate is when tuning DP-SGD with respect to standard SGD (with all the other conditions being the same). Definitively this work highlights how DP-tuning, while effectively working, requires to be handled in a slightly different way than a standard tuning.
@@
Li et al. (2021), building on the work of Papernot et al. (2019) on how to approach private-tuning, explore the effects of hyperparameters and fine-tuning objectives, identifying typical pitfalls of DP tuning when it’s treated as normal tuning. Their findings are particularly relevant as they suggest that, with proper execution, the performance of DP-tuned models can rival strong non-private benchmarks. They introduce a technique known as Ghost Clipping. Ghost Clipping is an important step in the landmark of private tuning because it greatly reduces memory usage, aligning the efficiency of private learning with that of non-private learning. Before that, DP-tuning was much slower and heavier than normal tuning to the point where it was almost not feasible (especially for large batch sizes, which are necessary for models like BERT) and even if it was, it would reach much lower performances. Ghost Clipping is an expansion of the concept described in (Lee & Kifer, 2020), it avoids the need to instantiate per-example gradients for each linear layer by handling one layer at a time and necessitating an additional backpropagation pass per batch. Extensive testing across various models—RoBERTa for natural language understanding (NLU), GPT2 for natural language generation (NLG), and both GPT-2 and DialoGPT for chit-chat dialogue generation—reveals that private models perform nearly as well as public ones, with only minor differences in perplexity and BLEU scores, showcasing competitive results. When compared to using vanilla Opacus, Ghost Clipping shows a 3x improvement in efficiency. While Ghost Clipping is already a big step in efficiency, Ding et al. (2024) point out that this method cannot be applied to standard Transformers because of non-plain feed-forward (and back-propagation) topology caused by embedding sharing that hinders Differential Privacy guarantees. The authors proposed Phantom Clipping, which is able to handle parameter sharing thanks to the manipulation of tensor operations. This ensures that the shared parameters’ gradients are computed and clipped without redundancy, hence preserving the Differential Privacy guarantees. They also show and solve another problem typical of DP-tuning: attention distraction. Attention distraction is a phenomenon that occurs when noise added in DP-SGD to maintain privacy distorts the self-attention mechanism, causing the model to inaccurately weigh parts of the input data. This leads to less effective learning as the model might focus on irrelevant details or overlook important information. To address this, the proposed "Re-Attention Mechanism" recalibrates attention scores by tracking and adjusting for the variance introduced by the noise. This adjustment ensures that attention scores more accurately reflect the true underlying data relationships despite the noise, thus maintaining the model’s performance and efficiency under differential privacy constraints.
@@
It is often possible that there is some overlap between pre-training data and fine-tuning data. While this is usually not an issue, in a private setting, where only fine-tuning is done privately and pre-training is standard, this kind of contamination would result in a data leakage that could expose sensitive data. inline]Furthermore, Tramèr et al. (2024) explained how this becomes even more of an issue as it is not clear if the reported results in these cases show an actual result for private learning or it is just a form of intra-context transfer learning. In this sense, the work done by Kurakin et al. (2024) is important as it addresses the potential data contamination between pre-training and fine-tuning datasets. To tackle this issue, they employ an LLM based on LaMDA (Thoppilan et al., 2022), initially applying deduplication techniques (Lee et al., 2022) to ensure fine-tuning data are excluded from the pre-training dataset. Following deduplication, they proceed with standard pre-training and subsequently conduct private fine-tuning using both full and Parameter Efficient Fine-Tuning (PEFT) methods. The model is then utilized to generate a synthetic dataset for a downstream sentiment analysis task, employing BERT and WordCNN for classification. Although the results are good, showing only minor performance drops for DP-tuning with respect to standard tuning, it is important to note the use of relatively high privacy budgets (e.g., $\epsilon$=10). Moreover, this approach is tested only on binary classification tasks, as opposed to the more complex multi-class classification challenges addressed by Li et al. (2021).
@@
Mattern et al. (2022a) introduce a variant of DP fine-tuning by applying it to GPT2 for generating private versions of IMDb and Amazon reviews datasets, which are then used to fine-tune models for classification. Following the methods of Schick & Schütze (2020), during the fine-tuning of GPT-2 they enhance sample generation fidelity by incorporating instructions based on the original dataset’s characteristics, such as sentiment or category, into templates. Since the datasets are made for classification, the data is labelled. The idea is to use these labels to steer the generation, also during training, by asking specifically for a sample corresponding to a label, like, for example, asking for a "positive" review. This approach integrates classification labels directly into the LLM’s fine-tuning prompts to prevent generating mismatched samples (like a positive review that reads like a negative one) and introduces a novel loss function that penalizes such mismatches. After training, the model samples a subset of the original training set, which is then utilized for further fine-tuning on binary classification tasks. For comparative analysis, the authors also fine-tune models using DP on the actual data. The main result is the demonstration that models trained on synthetic data generated via DP and then fine-tuned in a non-private manner outperform those trained directly on real data with DP. Additionally, the researchers assess data privacy by examining duplicates between original and synthetic datasets, noting a significant reduction, with duplicates completely eliminated in some instances. In this study, similarly to what was shown by Kurakin et al. (2024), the downstream task is binary classification, which is not particularly hard.

### v2

@@ 5 Solutions for Preserving Privacy in LLMs: Model
In this section, we explore the various contemporary methods employed to ensure privacy preservation within machine learning models. The focus is on four primary approaches that operate directly on the model. Firstly, we explore the concept of Differential Privacy applied during the training phase (§5.1). This method involves training a model with techniques designed to protect the privacy of the data used by injecting noise (for an overview on Differential Privacy, see §2.2). Secondly, we examine the application of Differential Privacy during the inference phase (§5.2), where privacy-preserving mechanisms are applied when the model is making predictions. Thirdly, we discuss federated learning with Differential Privacy (§5.3), which combines federated learning techniques with Differential Privacy to provide data safety in decentralized environments. Lastly, we discuss machine unlearning (§5.4), a technique that enables an existing model to forget or unlearn specific privacy-sensitive information, and Homomorphic Encryption (§5.5), a cryptographic method that allows computation on encrypted data without decrypting it. Each of these methods offers unique benefits and challenges in the pursuit of robust privacy protection in large language models.
@@ 5.1 Training Large Language Models with Differential Privacy
Training a Large Language Model (LLM) typically involves three sequential phases: pre-training on a broad collection of unannotated corpora (phase 1), fine-tuning on a specialized corpus using instruction-tuning procedures (phase 2), and final alignment according to human preferences (phase 3). We refer the reader to §2.1 for an in-depth overview of how these phases work. In this section, we propose methodologies that guarantee privacy by using differential privacy during the pre-training phase (phase 1 - §5.1.1) or apply it during the fine-tuning phase (phase 2 - §5.1.2). We also analyze methods that focus on Parameter Efficient Fine-Tuning techniques (§5.1.3) while leaving out of the scope of this survey works that focus on alignment via Reinforcement Learning and differential privacy since, to the best of our knowledge, there are very few works that just started exploring this aspect (Wu et al., 2024a).
Despite the significant concerns over privacy, most of today’s LLMs abstain from any privacy-specific measure for pre-training, relying instead on publicly available internet data for pre-training. While this data is public, it often contains sensitive information, posing risks that might harm individuals and undermine trust. Public data, thus, carries inherent risks yet remains a widely used resource. On the other hand, fine-tuning often requires using private datasets that entities prefer to keep confidential to protect sensitive information or maintain a competitive edge. We refer to "public training" when the training process involves using datasets that are available to the general public and to "private training" when the training process relies on restricted, confidential data that is not available to the public and must be carefully protected for privacy reasons. Here, Differential Privacy could offer a solution, minimizing risks associated with data memorization and safeguarding the privacy of individuals in these datasets.
We have already explained Differential Privacy (DP) in Section 2.2; here, we contextualize DP with its usage in modern large language models. Specifically, working with DP often involves deciding where to inject noise into the training process, typically by adding noise to the gradients, as shown in Step 2 of Figure 6. This approach generally involves using DP-SGD (Song et al., 2013), a variation of the classic SGD that clips gradient norms and then adds Gaussian noise to the average clipped gradient norm (see §2.2 for more details). It should be noted that, in principle, all the approaches involving DP-SGD are general, as DP-SGD is an optimizer that can be used with any network.
@@
Figure 6: Differential Privacy Applied on LMs. Differential privacy represents one of the main solutions used to preserve privacy when training a language model. We have already discussed Differential Privacy (DP) in Section 2.2; here, we contextualize DP with its usage in modern large language models that, most of the time, consist of adding noise in the training gradients (DP-SGD). On the right: Adding noise to the gradients of all the training parameters in a language model is often expensive. Parameter-efficient fine-tuning techniques can reduce the number of training parameters, thus lowering the cost of applying Differential Privacy to these gradients. Picture adapted from LoRA (Hu et al., 2022), a popular PEFT method.
@@ 5.1.1 Pre-training with Differential Privacy
Pre-training a Large Language Model from scratch is usually a very cost-intensive procedure that just a few companies and institutions in the world can afford. In addition to this, adding differential privacy to this pre-training phase further increases the difficulties since (i) adding noise increases the already exorbitant training cost, slowing the convergence time, and (ii) the training phase requires a vast collection of private data, which is often scarce and difficult to obtain. Thus, most models pre-trained from scratch with Differential Privacy are relatively small compared to current Large Language Model standards. Nevertheless, they offer both theoretical and practical insights into why fully private pre-training is considered a sub-optimal strategy. Instead, a mixed solution involving pre-training on public data is recommended. In this section, we will explain why. All the works discussed are also summarized in Table 4.
@@
It is crucial to acknowledge that the discussed works were capable of conducting full training from scratch largely because they utilized BERT (Devlin et al., 2019), which possesses a relatively modest 109M parameters, a figure significantly smaller than those of modern LLMs such as GPT-3 with 175B parameters (Brown et al., 2020). Currently, the standard practice, even in settings where data are not private, and there is no need to perform DP tuning, involves leveraging a pre-trained model and fine-tuning it for specific tasks. This approach not only reduces computational demands but also allows the model to harness a broad and diverse initial training dataset, enabling it to subsequently adapt more effectively to a private dataset and its specific domain, often enhancing performance. However, the benefits extend beyond performance improvements. Recent studies, including Kerrigan et al. (2020) that compare public pre-training with differentially private fine-tuning against entirely private DP training, demonstrate significantly better results with the former method. This provides empirical evidence that DP learning greatly benefits from an initial phase of public pre-training. Full training solely on a private dataset often results in high perplexity values, indicating a model that would likely be ineffective. In Kerrigan et al. (2020), the comparison of models fully trained privately versus those benefiting from public pre-training followed by private fine-tuning demonstrates that public pre-training is essential for enabling effective private learning. Additionally, since DP-SGD introduces noise during training, limiting this noise to the fine-tuning phase rather than the entire training process results in less overall noise and better performance. This improvement is due to the general (clean) knowledge acquired during the pre-training phase. Applying DP-SGD to an LLM pre-training (which is a huge part of the training) would make it extremely slow and add a lot of noise, which could hinder the final utility.
@@
Ganesh et al. (2023) present a study that compares the performance obtained via a public pre-training followed by private fine-tuning with a scenario where both stages are conducted in a differentially private manner. This paper emphasizes the theoretical underpinnings of Differential Privacy in training models, arguing that for a differentially private model to be effective on the task it is trained on, it is essential to first undergo pre-training on public data. The theoretical basis provided in the study explains that optimization involves two phases: initially, the model transitions from a random initialization towards a favourable basin, and subsequently, it addresses a convex optimization problem to find the local minimum within that basin. If attempted entirely within a private context, incorporating private data and noise addition, the initial stage would necessitate substantially more data to achieve comparable outcomes to those achievable with public data, as supported by experimental evidence. Moreover, given the typical scarcity of private data, it might be impractical or even impossible to gather sufficient data to effectively train a model privately from scratch, especially if the model is an LLM. Therefore, pre-training on a larger public dataset not only aids in overcoming these limitations but also enhances the model’s resilience to distribution shifts between the public and private datasets.
@@
Tramer & Boneh (2021) offer another perspective, arriving at a similar conclusion: public pre-training is crucial for achieving desirable outcomes in a private learning environment. Although this study is centred on image classification using CNNs, the findings reinforce that the advantages of public pre-training transcend various contexts, including this one. In their approach, the researchers enhance the performances of fully private training by incorporating handcrafted features, which yield improved results. Nevertheless, these outcomes still do not match those achieved through a combination of public pre-training and private fine-tuning. The collective insight from numerous studies is that a public pre-training phase is indispensable in private learning scenarios.
@@
While the works reviewed so far performed a "classical pre-training" (i.e., gaining knowledge from a large dataset), Yu et al. (2024) demonstrate that pre-training can also be conducted selectively, i.e., by focusing on a specific (smaller) subset of data that allows DP training to become more efficient. Their approach involves a novel technique of selecting a subset of the pre-training dataset, OpenWebText (Gokaslan & Cohen, 2019), whose distribution closely resembles the fine-tuning dataset’s distribution, Enron email dataset (Klimt & Yang, 2004). To filter the pre-training dataset, a classifier trained with Differential Privacy predicts whether a pre-training sample belongs to the distribution of the fine-tuning (private) dataset. Each sample in the pre-training dataset is then scored for how likely it is to belong to the fine-tuning dataset, then the top-k samples are used for public pre-training. This is followed by differentially private fine-tuning on the private dataset, with the final privacy loss taking into account both the classifier and the ultimate model. This method of dataset selection is designed to minimize distribution shifts, enabling smaller models to achieve better results. The results demonstrate strong performances, surpassing more complex baselines and thus enhancing efficiency to the point where this can be considered a model compression tool. However, a potential oversight of this method is the non-consideration of biases within the fine-tuning dataset, which could lead to the inadvertent transfer of these biases into the model through the selected pre-training samples. This is not as likely when the pre-training dataset is fully utilized. inline]While the learning process shows great benefit from a minimized distribution shift, this is not necessarily a good thing. As highlighted by Tramèr et al. (2024b), this can make the results less meaningful, as it becomes unclear if we are looking at an actual accomplishment for private learning or just another example of intra-domain transfer learning. They argue that most literature work use for pre-training and fine-tuning two datasets that are too similar, leading to these results only being valid if the distributions are close enough and making them invalid for fields where the publicly available data distribution is different than the private data distribution (which is often the case in the medical field). They also argue that this is fine if we are in the situation (called jokingly transfermania) where this is the norm, but if in the real world the distributions are usually different (privacyland) these results are not really meaningful.
@@ 5.1.2 Fine-tuning with Differential Privacy
Pre-training a large language model from scratch, with or without Differential Privacy, has become wildly expensive, given the sheer size of these models. Currently, the prevalent approach consists of using an already available LLM pre-trained checkpoint and fine-tuning it for specific downstream tasks, including applications of Differential Privacy (DP). In this section, we survey several different works that apply Differential Privacy on top of already pre-trained LMs. These works are also available in tabular format in Table 5. While the works in this section are all centered around LMs, they rarely adopt solutions specific only to the language domain, making these methods generalizable to other domains. Yue et al. (2023) show that using DP-SGD, it is possible to successfully train LMs for text generation. They apply DP-SGD to fine-tune various sizes of GPT2 using the Open Yelp Dataset. This is mainly done to generate synthetic reviews that cannot be traced to the actual reviewer. The generated reviews are then used to train classifiers to predict ratings and business categories. The quality of the synthetic text is assessed by comparing its distribution to the original dataset (through F1-score, Fréchet Inception Distance Heusel et al. (2017), and MAUVE Pillutla et al. (2021)) and the classifier performance against models trained on the original data. In this task, the largest GPT2 models produce data that more closely resemble the original data, probably because they are less influenced by DP. Experiments indicate that DP introduces a minimal performance drop, often less than 0.01%, within a privacy budget of $\epsilon=4$. Remarkably, the GPT2-Large model with DP nearly matches the performance of its non-private counterpart, with only a 0.008% difference in accuracy on downstream tasks. This leads to the notion that the pipeline is particularly suitable for LLMs, as it tends to scale well with model size. Privacy testing was performed following the method by Carlini et al. (2019): this consists of injecting canaries inside the training set in order to extract them later on, to prove that the model is capable (or not) of memorizing training data. When these canaries were injected into the training data, the model could not be prompted to reproduce them under DP conditions, even with a repetition rate of 100. In contrast, without DP, the canaries were reproduced 80% of the time, demonstrating effective privacy preservation. This work highlights the feasibility of DP fine-tuning.
@@
When we investigate DP learning and the used techniques, there is a constant comparison with standard learning, as it is the most widely known and adopted. Papernot et al. (2020) specifically focuses on the differences between private and public training, trying to shed some light on the best conditions for private training. They investigate architectures, parameter initializations and optimizers. We are not very interested in architectures as they do not consider transformers but only convolutional architectures and we already know from §5.1.1 that it’s best to start from parameters that were previously optimized through a public pre-training. The novelty that this research points out is that adaptive optimizers like Adam provide marginal gains (if any) in this context. Putting in direct comparison DP-SGD and DP-Adam, they find that while the latter converges faster initially, it eventually slows down to the point where DP-SGD can outperform it, reaching a lower loss. Another relevant difference they point out is how different a working learning rate is when tuning DP-SGD with respect to standard SGD (with all the other conditions being the same). Definitively this work highlights how DP-tuning, while effectively working, requires to be handled in a slightly different way than a standard tuning and, while this work only considers CNNs, there is no reason to believe that the findings would not apply to LLMs.
@@
Li et al. (2022), building on the work of Papernot et al. (2020) on how to approach private-tuning, explore the effects of hyperparameters and fine-tuning objectives, identifying typical pitfalls of DP tuning when it’s treated as normal tuning. Their findings are particularly relevant as they suggest that, with proper execution, the performance of DP-tuned models can rival strong non-private benchmarks. They introduce a technique known as Ghost Clipping. Ghost Clipping is an important step in the landmark of private tuning because it greatly reduces memory usage, aligning the efficiency of private learning with that of non-private learning. Before that, DP-tuning was much slower and heavier than normal tuning to the point where it was almost not feasible (especially for large batch sizes, which are necessary for models like BERT) and even if it was, it would reach much lower performances. Ghost Clipping is an expansion of the concept described in (Lee & Kifer, 2021), it avoids the need to instantiate per-example gradients for each linear layer by handling one layer at a time and necessitating an additional backpropagation pass per batch. Extensive testing across various models—RoBERTa for natural language understanding (NLU), GPT2 for natural language generation (NLG), and both GPT-2 and DialoGPT for chit-chat dialogue generation—reveals that private models perform nearly as well as public ones, with only minor differences in perplexity and BLEU scores, showcasing competitive results. When compared to using vanilla Opacus, Ghost Clipping shows a 3x improvement in efficiency. While Ghost Clipping is already a big step in efficiency, Ding et al. (2024) points out that this method cannot be applied to standard Transformers because of non-plain feed-forward (and back-propagation) topology caused by embedding sharing that hinders Differential Privacy guarantees. The authors proposed Phantom Clipping, which is able to handle parameter sharing thanks to the manipulation of tensor operations. This ensures that the shared parameters’ gradients are computed and clipped without redundancy, hence preserving the Differential Privacy guarantees. They also show and solve another problem typical of DP-tuning with transformers: attention distraction. Attention distraction is a phenomenon that occurs when noise added in DP-SGD to maintain privacy distorts the self-attention mechanism, typical of transformers networks, causing the model to inaccurately weight parts of the input data. This leads to less effective learning as the model might focus on irrelevant details or overlook important information. To address this, the proposed "Re-Attention Mechanism" recalibrates attention scores by tracking and adjusting for the variance introduced by the noise. This adjustment ensures that attention scores more accurately reflect the true underlying data relationships despite the noise, thus maintaining the model’s performance and efficiency under differential privacy constraints. It should be noted that these issues and fixes are general to transformers model and not specific to LLMs. However, they can greatly speed up fine-tuning for LLMs, which is quite important given their size.
@@
It is often possible that there is some overlap between pre-training data and fine-tuning data. While this is usually not an issue, in a private setting, where only fine-tuning is done privately and pre-training is standard, this kind of contamination would result in a data leakage that could expose sensitive data. inline]Furthermore, Tramèr et al. (2024b) explained how this becomes even more of an issue as it is not clear if the reported results in these cases show an actual result for private learning or it is just a form of intra-context transfer learning. In this sense, the work done by Kurakin et al. (2024) is important as it addresses the potential data contamination between pre-training and fine-tuning datasets. To tackle this issue, they employ an LLM based on LaMDA (Thoppilan et al., 2022), initially applying deduplication techniques (Lee et al., 2022) to ensure fine-tuning data are excluded from the pre-training dataset. Following deduplication, they proceed with standard pre-training and subsequently conduct private fine-tuning using both full and Parameter Efficient Fine-Tuning (PEFT) methods. The model is then utilized to generate a synthetic dataset for a downstream sentiment analysis task, employing BERT and WordCNN for classification. Although the results are good, showing only minor performance drops for DP-tuning with respect to standard tuning, it is important to note the use of relatively high privacy budgets (e.g., $\epsilon$=10). Moreover, this approach is tested only on binary classification tasks, as opposed to the more complex multi-class classification challenges addressed by Li et al. (2022).
@@
Mattern et al. (2022a) introduce a variant of DP fine-tuning by applying it to GPT2 for generating private versions of IMDb and Amazon reviews datasets, which are then used to fine-tune models for classification. Following the methods of Schick & Schütze (2020), during the fine-tuning of GPT-2, they enhance sample generation fidelity by incorporating instructions based on the original dataset’s characteristics, such as sentiment or category, into templates. Since the datasets are made for classification, the data is labelled. The idea is to use these labels to steer the generation, also during training, by asking specifically for a sample corresponding to a label, like, for example, asking for a "positive" review. This approach integrates classification labels directly into the LLM’s fine-tuning prompts to prevent generating mismatched samples (like a positive review that reads like a negative one) and introduces a novel loss function that penalizes such mismatches. Since this method takes advantage of prompts to steer the generation, it is specific to LLMs. After training, the model samples a subset of the original training set, which is then utilized for further fine-tuning on binary classification tasks. For comparative analysis, the authors also fine-tune models using DP on the actual data. The main result is the demonstration that models trained on synthetic data generated via DP and then fine-tuned in a non-private manner outperform those trained directly on real data with DP. Additionally, the researchers assess data privacy by examining duplicates between original and synthetic datasets, noting a significant reduction, with duplicates completely eliminated in some instances. In this study, similarly to what was shown by Kurakin et al. (2024), the downstream task is binary classification, which is not particularly hard.